# Learning Normals of Noisy Points by Local Gradient-Aware Surface Filtering

### v1

@@ 3 Method
[…]
@@ 3.1 Surface Fitting and Filtering
Figure 3: Left: computation of $f_{\theta}(\bm{q})\cdot\bar{\bm{n}}_{\bm{q}}$ and $\bar{\bm{n}}_{\bm{q}}\!=\!(\bm{n}_{\bm{q}}+\bm{n}_{\bm{q}^{\prime}})/||\bm{n}_{\bm{q}}+\bm{n}_{\bm{q}^{\prime}}||$. Gradients point to the positive side of the signed distance field. Right: computation of $\mathcal{H}(\bm{q})$ for specific noise and density using different neighborhood scales $K$.
@@ 3.3 Local Gradient Orientation of Intra-Level
Learning from multiple corrupt observations enhances performance on noisy data, but there remains room for improvement often overlooked by previous works [63, 32]. These approaches directly use $\bm{n}_{\bm{q}}$ as the normal of $\bm{q}$, neglecting the inaccuracies in gradients introduced by the local approximation in Eq. (8). We take this a step further by rethinking the generation of $\bm{q}^{\prime}$ from $\bm{q}$ and examining the geometric relationship of their gradients in the implicit field. For each query point $\bm{q}_{i}\!\in\!\bm{Q}$, we solve its normal as the sum of the normals at the two endpoints of the projection path, i.e., $\bar{\bm{n}}_{\bm{q}_{i}}\!=\!(\bm{n}_{\bm{q}_{i}}+\bm{n}_{\bm{q}_{i}^{\prime}})/||\bm{n}_{\bm{q}_{i}}+\bm{n}_{\bm{q}_{i}^{\prime}}||$, as illustrated in Fig. 3. The objective function then becomes the aggregation of errors at each neighborhood size scale:
$\mathcal{L}_{v}=\sum_{j=1}^{N_{K}}\frac{1}{N}\sum_{i=1}^{N}\left\|f_{\theta}(\bm{q}_{i})\cdot\frac{\bm{n}_{\bm{q}_{i}}+\bm{n}_{\bm{q}_{i}^{\prime}}}{||\bm{n}_{\bm{q}_{i}}+\bm{n}_{\bm{q}_{i}^{\prime}}||}-\mathcal{H}_{j}(\bm{q}_{i})\right\|.$ (10)
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China (62402401), the Sichuan Provincial Natural Science Foundation of China (2025ZNSFSC1462) and the Fundamental Research Funds for the Central Universities.
@@ Appendix A More Experimental Results
Figure 13: Visual comparison of point cloud denoising. Our method is unsupervised, whereas all baseline methods rely on supervised training. The input consists of 50K-resolution shapes with 3% Gaussian noise from the PointCleanNet dataset. Point colors represent P2M distance error (×$\times 10^{-4}$).
@@ Appendix B Discussion
Hyperparameter tuning in loss. Compared to prior works, we enhance existing losses from a new perspective and introduce novel richer constraints to boost performance. To ensure each loss term contributes meaningfully and no term is orders of magnitude larger or smaller, we let terms with larger raw values receive smaller weight factors, and vice versa. The weights $\lambda_{1}$ and $\lambda_{2}$ are chosen via small grid searches on a held-out subset. We find that varying each weight by ±$\pm 50\%$ changes the average RMSE by $<3\%$, showing low sensitivity. For $\lambda_{1}$, we recommend trying $\{0.05,0.1,0.3\}$ to choose the value that minimizes normal RMSE. $\lambda_{2}$ are less sensitive and can remain the defaults unless targeting extremely noisy or highly detailed data. In $\mathcal{L}_{sd}$, a 10×10$\times$10 × factor on the signed distances of surface points forces them to be located on the zero level set and balances surface fitting and denoising. We will release code with default settings for transparency.

### v2

@@ 3 Method
Preliminary. Implicit representation approaches usually denote surfaces as the level sets of implicit function, i.e., $\bm{S}_{d}\!=\!\left\{\bm{x}\in\mathbb{R}^{3}~|~f_{\theta}(\bm{x})\!=\!d\right\}$, where $f_{\theta}\colon\mathbb{R}^{3}\!\rightarrow\!\mathbb{R}$ is implemented as a neural network with parameter $\theta$. The implicit function can be learned by overfitting the neural network on individual point clouds. If the function $f_{\theta}$ is correctly defined by a signed distance field inferred from points, the normal of a point $\bm{p}$ in this implicit field can be obtained by $\bm{n}_{\bm{p}}\!=\!\nabla f_{\theta}(\bm{p})/\left\|\nabla f_{\theta}(\bm{p})\right\|$, where $\left\|\cdot\right\|$ means the Euclidean $L^{2}$-norm and $\nabla f_{\theta}(\bm{p})$ denotes the gradient at $\bm{p}$. Specifically, the zero level set $f_{\theta}(\bm{x})\!=\!0$ is usually extracted as the object or scene surface $\bm{S}$. Random points on a level set have specific signed distances, such as $f_{\theta}(\bm{x})\!<\!0$ for outside and $f_{\theta}(\bm{x})\!>\!0$ for inside. The gradients on a specific iso-surface should have uniform orientations. In this work, we aim to apply surface filtering to project noisy points onto the underlying surface defined by the zero level set without supervision of ground truth labels or clean points. We use the signed distances and normals of noisy points to define the projection path and incorporate rules of the local field. An overview of our method is provided in Fig. 1.
@@ 3.1 Surface Fitting and Filtering
[…]
@@ 3.3 Local Gradient Orientation of Intra-Level
[…]
@@ Acknowledgements
This work was supported by the National Natural Science Foundation of China (62402401), the Sichuan Provincial Natural Science Foundation of China (2025ZNSFSC1462) and the Fundamental Research Funds for the Central Universities (2682025CX109).
@@ Appendix A More Experimental Results
Figure 13: Visual comparison of point cloud denoising. Our method is unsupervised, whereas all baseline methods rely on supervised training. The input consists of 50K-resolution shapes with 3% Gaussian noise from the PointCleanNet dataset. Point colors represent P2M distance error ($\times 10^{-4}$).
@@ Appendix B Discussion
Hyperparameter tuning in loss. Compared to prior works, we enhance existing losses from a new perspective and introduce novel richer constraints to boost performance. To ensure each loss term contributes meaningfully and no term is orders of magnitude larger or smaller, we let terms with larger raw values receive smaller weight factors, and vice versa. The weights $\lambda_{1}$ and $\lambda_{2}$ are chosen via small grid searches on a held-out subset. We find that varying each weight by $\pm 50\%$ changes the average RMSE by $3\%$, showing low sensitivity. For $\lambda_{1}$, we recommend trying $\{0.05,0.1,0.3\}$ to choose the value that minimizes normal RMSE. $\lambda_{2}$ are less sensitive and can remain the defaults unless targeting extremely noisy or highly detailed data. In $\mathcal{L}_{sd}$, a $10\times$ factor on the signed distances of surface points forces them to be located on the zero level set and balances surface fitting and denoising. We will release code with default settings for transparency.